A girl is standing at the edge of a cliff 100. m above the ground. She reaches out over the edge of the cliff and throws a rock straight upward


with a speed 8.00 m/s.


a) How long does it take the rock to hit the ground?


b) What is the speed of the rock the instant before it hits the ground?



I think the answers are:


a) 6.4s


b) 45.0m

1
Expert's answer
2021-11-07T19:24:50-0500

"h'=\frac{v^2}{2g}=3.3~m,"

"H=h+h'=103.3~m,"

a)

"t=t_1+t_2=\sqrt{\frac{2h'}g}+\sqrt{\frac{2H}g}=5.4~s,"

b)

"u=\sqrt{2gH}=45~\frac ms."